OpenStudy (anonymous):

Under certain conditions, neon (Ne) gas diffuses at a rate of 7.0 centimeters per second. Under the same conditions, an unknown gas diffuses at a rate of 4.9 centimeters per second. What is the approximate molar mass of the unknown gas?

OpenStudy (anonymous):

6.5 g/mol 11 g/mol 20 g/mol 41 g/mol

OpenStudy (aaronq):

use grahams law of diffusion: \(\large \dfrac{rate_1}{rate_2}=\sqrt{\dfrac{M_2}{M_1}}\) M=molar mass

OpenStudy (anonymous):

so 7.0/4.9 = ?/20.17

OpenStudy (anonymous):

@aaronq

OpenStudy (aaronq):

yep, but dont forget the squared root

OpenStudy (aaronq):

do you know how to solve for \(M_2\)?

OpenStudy (anonymous):

Nope

OpenStudy (aaronq):

okay, it's just algebra \(\large\dfrac{rate_1}{rate_2}=\sqrt{\dfrac{M_2}{M_1}}\) square both sides \(\bigg(\large\dfrac{rate_1}{rate_2}\bigg)^2=\bigg(\sqrt{\dfrac{M_2}{M_1}}\bigg)^2\) the squared root sign cancels by squaring it \(\bigg(\large\dfrac{rate_1}{rate_2}\bigg)^2=\dfrac{M_2}{M_1}\) move \(M_1\) over \(\large M_2=M_1*\bigg(\large\dfrac{rate_1}{rate_2}\bigg)^2\)

OpenStudy (anonymous):

M2 = 41.14?

OpenStudy (aaronq):

yup

OpenStudy (anonymous):

So I got 1.42

OpenStudy (anonymous):

Which isn't an answer, so is there something I do next?

OpenStudy (aaronq):

your answer was \(M_2\) = 41

OpenStudy (anonymous):

Oh

OpenStudy (anonymous):

Thank you!